OpenStudy (anonymous):

Help will givemedal Factor: 36t^2-81y^2

OpenStudy (anonymous):

what might the question be?

OpenStudy (anonymous):

ooh factor difference of two squares \[a^2-b^2=(a+b)(a-b)\] with \[a=6t, b=9y\]

OpenStudy (anonymous):

is that how it is ?

OpenStudy (ybarrap):

Yes!

OpenStudy (anonymous):

and from there /

OpenStudy (ybarrap):

Just plug in a and b and you're done!

OpenStudy (ybarrap):

$$ a^2-b^2=(a+b)(a-b)\\ a=6t, b=9y\\ (6t)^2-(9y)^2=? $$

OpenStudy (anonymous):

(2t+3y) (2t-3y) ?

OpenStudy (ybarrap):

No. $$ a=6t, b=9y $$ Then just plug in a and b here: $$ (a+b)(a-b) $$

OpenStudy (anonymous):

so, (6t+9y) (6t-9y) ?